The equation of motion of projectile are given by x=36t and 2y=96t-9.8t×t
 

Grade:11

2 Answers

Arun
5 years ago
Dear Devansh
 
question posted is incomplete. Please check and repost this. You canalso attach an image.
 
I will be happy to help ypu.
Khimraj
5 years ago

Comparing x = 36t with x u cos t, u cos  θ =36 

2y =  96t -9.8t2

which imlpies  y =48t -4.9t2

comparing y with (usinθ )t – gt2/2

usinθ  = 48

usinθ/ucosθ = 48/36 = 4/3

tanθ  = 4/3

θ = tan-1(4/3)

Hope it clears.................